# GPT-OSS-20B: A Comprehensive Deployment-Centric Analysis of OpenAI’s Open-Weight Mixture of Experts Model

### v1

@@ 1 Introduction
The landscape of open-weight large language models has been reshaped by Mixture-of-Experts (MoE) architectures, which activate only a subset of parameters per token and thereby reduce inference cost relative to dense models of similar total size [1]. In this study, we evaluate an open-weight MoE checkpoint, GPT-OSS-20B, which has approximately 20.9B total parameters but only $\sim$3.61B active at inference (17.3% active fraction) [3]. The model was released under a permissive open-source license [2].
@@ 1.1 Motivation
The deployment of large language models in production environments requires careful consideration of multiple factors beyond accuracy, including latency, throughput, memory efficiency, and energy consumption. Traditional dense models, while achieving high accuracy, often require significant computational resources that may not be available in resource-constrained environments. Mixture-of-Experts (MoE) architectures offer a promising alternative by activating only a subset of parameters during inference, potentially reducing resource requirements while maintaining competitive performance.
@@ 1.2 Contributions
    Unified, reproducible deployment benchmarking. A single-GPU (H100, bf16) harness for latency/throughput (true TTFT, p50/p95, TPOT from full decode), memory (peak VRAM with PKV held), and energy (tokens per Joule, J/1K via stabilized nvidia-smi sampling).
@@ 2.1 Mixture of Experts Models
Mixture of Experts (MoE) models have emerged as a promising approach to scale language models efficiently. The key insight is that not all parameters need to be active during inference, allowing for larger models with manageable computational requirements. Foundational works include the sparsely-gated MoE layer [1], GShard [9], and the Switch Transformer [10], which demonstrated that MoE can achieve strong performance while significantly reducing active parameter counts.
@@ 2.2 Deployment-Centric Evaluation
While most evaluations focus on accuracy metrics, deployment characteristics are crucial for real-world applications. Efficiency aspects such as latency, throughput, and energy consumption have been highlighted in recent surveys of resource-efficient LLMs [4] as well as standardized benchmarks such as MLPerf Inference [5]. However, systematic comparisons of MoE versus dense models under deployment constraints remain limited.
@@ 2.3 Open-Weight Model Evaluation
Recent evaluations of open-weight models such as Mistral [6], Qwen [7], and Yi [8] have emphasized accuracy benchmarks. Deployment-centric trade-offs, particularly between MoE and dense models, remain underexplored.
@@ 3.1 Experimental Setup
Hardware. All measurements were taken on a single NVIDIA H100 GPU (bf16), with no sharding or offload. We pin to one device, clear caches between runs, and hold the persistent KV cache (PKV) in memory during measurement.
@@ 3.4 Peak Memory Measurement
  • We read max_memory_allocated() as Peak VRAM, which captures KV cache plus transient kernels.
@@ 3.5 Energy Measurement
  • Power from nvidia-smi is approximate; using identical sampling and run structure across models makes the comparisons reliable.
@@ 3.6 Active Parameter Efficiency (APE)
APE contextualizes performance by the fraction of parameters active at inference:
@@
TPOT/GB $=\frac{\text{TPOT}}{\text{PeakMem (GB)}}\quad\text{(auxiliary, not an APE term)}.$
@@ 3.7 Ablation Protocols
  • Decoding: greedy vs. sampling (top-$p$, top-$k$; temperature sweeps).
@@
  • Serving stack: direct transformers runs; comparisons with alternative servers (e.g., vLLM) require separate deployment and are not included in the core numbers.
@@ 4.2 Memory Analysis
Table 5: Peak VRAM at context =2,048=2{,}048 and decode =64=64 tokens (allocator peak with PKV alive). Lower is better. $\Delta$ vs. GPT-OSS” is an absolute MB gap; “% less vs. GPT-OSS” uses the baseline’s memory as the denominator.
Model Peak VRAM (MB) $\Delta$ vs. GPT-OSS (MB) % less vs. GPT-OSS
@@
At 2,048 context, GPT-OSS-20B uses 31.71% and 34.60% less peak VRAM than Qwen3-32B and Yi-34B, respectively (Table 5), calculated relative to each model’s peak VRAM. In absolute terms, this corresponds to reductions of $\sim$20.2 GB and $\sim$23.0 GB versus the dense 30–34B baselines on a single H100. These savings stem from GPT-OSS-20B’s MoE architecture, which requires less baseline memory (39.9 GB) compared to Qwen3-32B (62.5 GB) and Yi-34B (65.6 GB), despite a larger KV cache footprint, measured with identical post-template token counts.
Caveats.
Allocator peaks include transient workspace from fused kernels and can vary slightly across library versions. We mitigate this by (i) a dedicated warm-up, (ii) resetting peak counters prior to the measured run, and (iii) holding PKV alive through the readout. Results reflect a single-GPU (H100, bf16) configuration with no sharding or offload.
4.3 Energy Analysis
We estimate energy using nvidia-smi power samples on repeated short runs (one warm-up), then compute throughput (TPOT), tokens per watt, and Joules per 1K generated tokens from the measured wall time. We report medians over repeats and enforce the same post-template context as in latency/memory.
@@ 4.3 Energy Analysis
Relative to Qwen3-32B, GPT-OSS-20B delivers +31.8% higher TPOT, +34.7% higher tokens/W, and -25.8% lower J/1K generated tokens (Table 6). Versus Yi-34B, it shows +18.9% TPOT, +37.9% higher tokens/W, and -27.5% lower J/1K.
@@ 4.5 Ablation Studies (GPT-OSS-20B)
vLLM Not run (server out of scope)
@@ 4.6 Safety and Governance (Qualitative)
Provide authoritative citations for license/policy/governance, augment with quantitative safety evaluation (harmlessness and jailbreak suites), and document any deployment-time filters or guardrails alongside model results.
@@ 5 Conclusion
At 2048-token contexts, GPT-OSS-20B delivers higher throughput, lower energy per 1,000 generated tokens, higher tokens per Joule, and substantially lower peak memory than dense baselines Qwen3-32B and Yi-34B in our single-GPU (H100, bf16) setup, though with higher TTFT due to its MoE architecture. APE complements raw metrics by normalizing for active parameters, highlighting GPT-OSS-20B’s efficiency per active parameter. We release code and consolidated results to facilitate replication.

### v2

@@ 1 Introduction
The landscape of open-weight large language models has been reshaped by Mixture-of-Experts (MoE) architectures, which activate only a subset of parameters per token and thereby reduce inference cost relative to dense models of similar total size shazeer2017outrageously . In this study, we evaluate an open-weight MoE checkpoint, GPT-OSS-20B, which has approximately 20.9B total parameters but only $\sim$3.61B active at inference (17.3% active fraction) openai2025gptossmodelcard . The model was released under a permissive open-source license openai2025gptoss
@@ 1.1 Motivation
The deployment of large language models in production environments requires careful consideration of multiple factors beyond accuracy, including latency, throughput, memory efficiency, and energy consumption. Traditional dense models, while achieving high accuracy, often require significant computational resources that may not be available in resource-constrained environments. Mixture-of-Experts (MoE) architectures offer a promising alternative by activating only a subset of parameters during inference, potentially reducing resource requirements while maintaining competitive performance jiang2024mixtral
@@ 1.2 Contributions
    Unified, reproducible deployment benchmarking. A single-GPU (H100, bf16) harness for latency/throughput (true TTFT, p50/p95, TPOT from full decode), memory (peak VRAM with PKV held), and energy (tokens per Joule, J/1K via stabilized nvidia-smi sampling). nvsmi_accuracy_issues_2024
@@ 2.1 Mixture of Experts Models
Mixture of Experts (MoE) models have emerged as a promising approach to scale language models efficiently. The key insight is that not all parameters need to be active during inference, allowing for larger models with manageable computational requirements. Foundational works include the sparsely-gated MoE layer shazeer2017outrageously , GShard lepikhin2020gshard , ST-MoE stmoe and the Switch Transformer fedus2021switch , which demonstrated that MoE can achieve strong performance while significantly reducing active parameter counts.
Recent advancements have built on these foundations, with models like Mixtral 8x7B jiang2024mixtral introducing sparse MoE layers for improved efficiency in open-weight settings. Similarly, DeepSeek-MoE dai2024deepseekmoe explores routing strategies to balance load across experts, achieving up to 70% reduction in active parameters during inference. Grok-1 xai2024grok1 , an open-source MoE model with 314B parameters, further demonstrates scaling to massive sizes while maintaining low inference costs through optimized expert selection. Optimized frameworks like DeepSpeed-MoE rajbhandari2022deepspeed support efficient training and inference for such models. A comprehensive survey by Cai et al. cai2024survey provides a taxonomy of MoE designs, covering algorithmic aspects like gating functions and expert architectures, as well as systemic considerations such as computation, communication, and storage optimizations.
2.2 Efficiency and Compression Techniques for MoE Models
Beyond architectural innovations, recent research has focused on enhancing the efficiency of MoE models through compression and optimization techniques. Su et al. su2025unveiling identify “Super Experts” in MoE LLMs—a small subset of experts critical for model performance, characterized by extreme activation patterns. Pruning these leads to significant degradation, particularly in mathematical reasoning, highlighting the need to preserve them during compression.
Huang et al. huang2025mixture propose the Mixture Compressor (MC), a training-free method combining mixed-precision quantization and dynamic pruning. It achieves substantial compression (e.g., 76.6% at 2.54 bits) with minimal accuracy loss by considering expert importance and token criticality, further reducing activated parameters during inference. Complementary post-training quantization methods apply to MoE experts and routers, including LLM.int8 dettmers2022llmint8 , GPTQ frantar2022gptq , SmoothQuant xiao2022smoothquant , AWQ lin2023awq , and QServe lin2024qserve .
These works underscore the importance of targeted efficiency improvements in MoE models, balancing model size with deployment feasibility.
@@ 2.3 Deployment-Centric Evaluation
While most evaluations focus on accuracy metrics, deployment characteristics are crucial for real-world applications. Efficiency aspects such as latency, throughput, and energy consumption have been highlighted in recent surveys of resource-efficient LLMs liu2023survey as well as standardized benchmarks such as MLPerf Inference reddi2020mlperf . However, systematic comparisons of MoE versus dense models under deployment constraints remain limited. Distributed serving systems like Orca yu2022orca address scalability for large models, as demonstrated in recent MLPerf results using TensorRT-LLM.
Inference optimizations like FlashAttention dao2022flashattention ; dao2023flashattention2 ; shah2024flashattention3 and PagedAttention kwon2023efficient have been proposed to reduce memory and latency in transformer-based models, particularly for long contexts. Studies on energy efficiency, such as those in poddar2025sustainable ; patel2024asplos ; patel2024isca ; wilkins2024hotcarbon , analyze power consumption in LLM serving, emphasizing the need for hardware-aware metrics like tokens per Joule. Recent libraries like FlashInfer flashinfer2025lib provide efficient attention engines tailored for LLM serving.
A survey by Chang et al. chang2024survey reviews LLM evaluation across tasks, methods, and benchmarks, stressing the importance of holistic assessment including efficiency. Saleh et al. saleh2025evaluating provide a systematic review of LLM efficiency, applications, and future directions, analyzing models like GPT-3 and Codex in terms of hardware setups, parameters, and performance metrics.
@@ 2.4 Open-Weight Model Evaluation
Recent evaluations of open-weight models such as Mistral jiang2023mistral , Qwen team2024qwen , and Yi ai2024yi have emphasized accuracy benchmarks. Deployment-centric trade-offs, particularly between MoE and dense models, remain underexplored. For instance, evaluations of Llama models touvron2023llama highlight scaling laws but often overlook single-GPU energy and memory profiles in production-like settings. Such evaluations often align with scaling laws from works like Chinchilla hoffmann2022chinchilla .
2.5 Energy Efficiency and Sustainability in LLM Inference
The increasing computational demands of large language models (LLMs) have spurred research into energy-efficient inference strategies, particularly for resource-constrained environments. Maliakel (2025) investigates energy-performance trade-offs in LLM inference, demonstrating that dynamic voltage and frequency scaling (DVFS) can reduce power consumption by up to 30% with minimal latency increases across various tasks (maliakel2025, ). Fernandez et al. (2025) explore quantization and pruning techniques, achieving up to 73% energy reduction in NLP tasks by optimizing model compression without significant performance degradation (fernandez2025, ). Poddar et al. (2025) (poddar2025sustainable, ) benchmark inference energy across diverse LLMs, identifying hardware-specific factors that influence energy profiles and advocating for standardized tokens-per-Joule metrics. Dauner and Socher (2025) quantify the environmental impact of LLM interactions, emphasizing CO2 emissions and proposing energy-aware metrics to guide sustainable deployments (dauner2025, ). Saleh et al. (2025) saleh2025evaluating provide a comprehensive review of LLM efficiency, highlighting hardware-aware optimizations and future directions for reducing energy footprints in production settings. These studies align with the need for metrics like Active Parameter Efficiency (APE), which normalize energy consumption by active parameters, to enhance the sustainability of MoE model deployments.
@@ 3.1 Experimental Setup
Hardware. All measurements were taken on a single NVIDIA H100 GPU (bf16), with no sharding or offload. We pin to one device, clear caches between runs, and hold the persistent KV cache (PKV) in memory during measurement kwon2023efficient .
@@
Units.
Unless otherwise stated, MB/GB are decimal (1 MB=106 bytes, 1 GB=109 bytes). We use GiB (1 GiB=230 bytes) only when explicitly labeled.
@@ 3.4 Peak Memory Measurement
  • We read torch.cuda.max_memory_allocated() as Peak VRAM, which captures KV cache plus transient kernels.
@@ 3.5 Energy Measurement
  • Power from nvidia-smi is approximate; using identical sampling and run structure across models makes the comparisons reliable. Alternative tools for energy tracking include EIT henderson2020eit and eco2AI eco2ai_2022 with benchmarking methodology reinforced by Pope et al. pope2023benchmarking .
@@ 3.6 Active Parameter Efficiency (APE)
While prior studies have examined scaling efficiency in terms of total parameters hoffmann2022chinchilla ; du2022glam , they do not account for the sparsity properties of Mixture-of-Experts models, where only a fraction of weights are active at inference. To address this, we introduce Active Parameter Efficiency (APE) as a normalization lens that contextualizes deployment metrics by the number of parameters actually used during inference. APE provides a per-active-parameter view of throughput, latency, and energy, enabling apples-to-apples comparisons between dense and sparse models:
@@
TPOT/GB $=\frac{\text{TPOT}}{\text{PeakMem (GB)}}\quad\text{(auxiliary, decimal GB)}.$
This formulation highlights how much performance is delivered per active parameter (or per gigabyte of peak memory), complementing raw deployment metrics with a sparsity-aware efficiency perspective.
@@ 3.7 Ablation Protocols
  • Decoding: greedy vs. sampling (top-$p$, top-$k$; temperature sweeps). We do not explore advanced decoding techniques like speculative decoding leviathan2022specdec ; liu2024specdec , which could further accelerate inference.
@@
  • Serving stack: direct transformers runs; comparisons with alternative servers (e.g., vLLM) require separate deployment and are not included in the core numbers. kwon2023efficient Other optimized backends include NVIDIA’s TensorRT-LLMnvidia2024tensorrtllm for high-performance inference for CPU/GPU portability, with optimizations such as CUDA graphs gtc2023_cudagraphs .
@@ 4.2 Memory Analysis
Table 5: Peak VRAM at context =2,048=2{,}048 and decode =64=64 tokens (allocator peak with PKV alive). Lower is better. $\Delta$ vs. GPT-OSS” is an absolute MB gap; “% less vs. Qwen/Yi” uses the baseline’s memory as the denominator.
Model Peak VRAM (MB) $\Delta$ vs. GPT-OSS (MB) % less vs. Qwen/Yi
@@
At 2,048 context, GPT-OSS-20B uses 31.71% and 34.60% less peak VRAM than Qwen3-32B and Yi-34B, respectively (Table 5), calculated relative to each model’s peak VRAM. In absolute terms, this corresponds to reductions of $\sim$20.2 GB and $\sim$23.0 GB versus the dense 30–34B baselines on a single H100. These savings stem from GPT-OSS-20B’s MoE architecture, which requires less baseline memory ($\approx$41.8 GB, decimal) compared to Qwen3-32B (62.5 GB) and Yi-34B (65.6 GB), despite a larger KV cache footprint, measured with identical post-template token counts. Here, baseline memory means VRAM after weights load but before any input, excluding KV and transient kernels. All memory results are allocator peaks with PKV held, reported as medians over repeats at matched contexts.
@@
Relative to Qwen3-32B, GPT-OSS-20B delivers +31.8% higher TPOT, +34.2% higher tokens/W, and $25.8\%$ lower J/1K generated tokens (Table 6). Versus Yi-34B, it shows +18.9% TPOT, +37.8% higher tokens/W, and $27.5\%$ lower J/1K.
@@ 4.4 Ablation Studies (GPT-OSS-20B)
vLLM kwon2023efficient . Not run (server out of scope)
@@ 4.5 Safety and Governance (Qualitative)
Future work on deployment-focused evaluations should incorporate authoritative, peer-reviewed studies on license, policy, and governance considerations weidinger2021ethical ; bommasani2021foundation . In addition to qualitative documentation, systematic safety assessments—such as harmlessness and jailbreak benchmarks—are necessary to provide a more complete picture of model behavior. Finally, deployment reports should explicitly describe any runtime guardrails or filtering mechanisms applied during serving, ensuring transparency and reproducibility.
@@ 5 Conclusion
At 2048-token contexts, GPT-OSS-20B delivers higher throughput, lower energy per 1,000 generated tokens, higher tokens per Joule, and substantially lower peak memory than dense baselines Qwen3-32B and Yi-34B in our single-GPU (H100, bf16) setup, though with higher TTFT due to its MoE architecture. APE complements raw metrics by normalizing for active parameters, highlighting GPT-OSS-20B’s efficiency per active parameter. These results suggest MoE models are more viable for single-GPU deployment in production settings than dense models of similar scale.